OpenStudy (anonymous):

What is the slope of the line that goes through (6, 5) and (3, 1)?

OpenStudy (anonymous):

for any two point x1,y1 and (x2,y2) slope is \[m=\frac{ (y _{1} - y _{2}) }{ (x _{1} -x _{2} ) }\]

OpenStudy (anonymous):

so m=5-1/6-3=4/3

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

its ok and please Memories the formula for next time.